# Deep Learning and Machine Learning - Python Data Structures and Mathematics Fundamental: From Theory to Practice

### v1

@@ 15.2 Frobenius Norm
The Frobenius norm is a matrix norm equivalent to the L2 norm for vectors[227]. It is defined as the square root of the sum of the absolute squares of the matrix elements:
@@ 15.3 Cosine Similarity
Cosine similarity measures the cosine of the angle between two vectors[180]. It is often used in text analysis and other fields where the direction of vectors is more important than their magnitude.
@@
cosine $\text{cosine similarity}=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$
@@ 15.4 Euclidean Distance
Euclidean distance is the straight-line distance between two points in Euclidean space[63]. For two vectors $\mathbf{a}$ and $\mathbf{b}$, it is defined as:
@@
Automatic differentiation (AD) is a powerful technique used in many machine learning frameworks, including PyTorch and TensorFlow, to compute gradients efficiently and accurately[16]. Unlike symbolic differentiation, which can produce complex expressions, or numerical differentiation, which can suffer from precision issues, automatic differentiation computes derivatives systematically using the chain rule. This chapter introduces the concept of automatic differentiation and demonstrates how gradients can be computed in popular machine learning libraries like PyTorch and TensorFlow.
@@ 16.1 Introduction to Automatic Differentiation
Reverse-mode AD is particularly useful in deep learning, where we often need to compute the gradient of a loss function with respect to model parameters[89].
@@ 16.4 Jacobian and Hessian Computation
In some cases, we may need to compute higher-order derivatives, such as the Jacobian matrix or Hessian matrix[140]. The Jacobian represents the matrix of first-order partial derivatives for a vector-valued function, while the Hessian is the matrix of second-order partial derivatives.
@@ 16.5 Summary
In this chapter, we will introduce the basic concepts behind optimization in deep learning, starting with Gradient Descent, which is the foundation of many advanced optimization techniques. We will then explore Stochastic Gradient Descent (SGD)[232], momentum-based optimization, and adaptive optimization methods like Adagrad[76], RMSprop[266], Adam[148], and AdamW[176].
@@ 17.1 Gradient Descent
Gradient Descent is one of the most fundamental optimization algorithms used in deep learning. The basic idea is to minimize a loss function (also called the objective function) by moving in the direction of the negative gradient, which points towards the steepest descent[17].
@@ 17.2 Stochastic Gradient Descent (SGD)
While Gradient Descent computes the gradient using the entire dataset, this can be computationally expensive for large datasets. Stochastic Gradient Descent (SGD) solves this problem by approximating the gradient using a small batch of the data, or even a single data point[199].
@@ 17.3 Momentum-based Optimization
Momentum-based optimization improves the convergence speed of Gradient Descent by adding momentum to the update rule[64]. This allows the optimizer to continue moving in the same direction if the gradient consistently points in the same direction, avoiding oscillations and speeding up convergence.
@@ 17.4 Adaptive Optimization Methods
Adaptive optimization methods automatically adjust the learning rate during training, which can significantly improve convergence[173]. These methods include Adagrad, RMSprop, Adam, and AdamW, each of which modifies the learning rate based on the gradients.
@@ 17.5 Learning Rate Schedules
The learning rate is one of the most important hyperparameters in deep learning. It controls how much the model’s weights are adjusted with respect to the loss gradient during training. Choosing the correct learning rate can significantly impact model performance and training time. However, the optimal learning rate often changes throughout training. This is where learning rate schedules come into play[66].
A learning rate schedule adjusts the learning rate dynamically during training, helping the model converge faster and avoid getting stuck in local minima. There are several common strategies for learning rate scheduling, including step decay, exponential decay, and warm restarts[177].
@@ 17.5.1 Step Decay
Step decay is one of the simplest learning rate schedules. In step decay, the learning rate is reduced by a constant factor at predetermined intervals (epochs)[282]. The idea is that a larger learning rate is beneficial at the start of training, but as the model approaches convergence, reducing the learning rate allows finer adjustments to the weights.
@@ 17.5.2 Exponential Decay
Exponential decay is another commonly used learning rate schedule. Instead of reducing the learning rate in steps, exponential decay reduces the learning rate continuously over time according to an exponential function[166]. This allows for a more gradual and smoother reduction in the learning rate.
@@ 18.1 Batch Normalization in Training
Batch normalization is a technique that helps to stabilize and accelerate the training of deep neural networks by normalizing the inputs to each layer[130]. This process ensures that the input distribution to each layer remains stable during training, making the network less sensitive to the initialization of parameters and improving convergence.
@@ 18.2 Gradient Clipping
Gradient clipping is a technique used to prevent the problem of exploding gradients during the training of deep neural networks[210, 174]. When gradients become excessively large, they can cause updates to the network weights that are too drastic, leading to unstable training or even causing the model to diverge. Gradient clipping limits the size of the gradients by setting a maximum threshold.
@@ 18.2.1 How Does Gradient Clipping Work?
Given a gradient $g$ and a threshold $t$, if $||g||>t$, we rescale the gradient:
$g_{\text{clipped}}=g\cdot\frac{t}{||g||}$
@@ 18.3 Second-order Optimization Methods
Second-order optimization methods are more sophisticated than standard first-order methods like gradient descent[15]. These methods take into account not only the gradient (first derivative) but also the curvature of the objective function (second derivative). This can lead to faster convergence in some cases, especially when the objective function has complex curvature.
@@ 18.3.1 Newton’s Method
Newton’s Method is a second-order optimization technique that uses both the gradient and the Hessian matrix (the matrix of second-order partial derivatives) to update parameters[216]. The basic update rule in Newton’s Method is:
@@ 18.3.2 Quasi-Newton Methods
Quasi-Newton methods are a family of optimization algorithms that approximate the Hessian matrix rather than computing it explicitly[87]. This makes them more computationally efficient than full Newton’s Method, especially for high-dimensional problems. One of the most popular Quasi-Newton methods is the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm[101].
@@ 20.2 Basic Gradient Computation Problems
Gradient computations are essential for optimization in deep learning, as they enable the backpropagation algorithm to update network weights[17]. In this section, we will practice calculating gradients manually and using automatic differentiation tools like those in NumPy.
@@ 20.4 Real-World Linear Algebra Applications in Deep Learning
In deep learning, linear algebra concepts like matrix multiplication, matrix inversion, and eigenvalue decomposition are used extensively[4]. Here are some real-world applications:
@@ 22.2 Sources of Errors in Computation
  • Round-off Error[72]: This occurs due to the limited precision of floating-point arithmetic used by computers. For example, irrational numbers like $\pi$ and square roots of non-perfect squares cannot be represented exactly.
  • Truncation Error[256]: This occurs when an infinite process is approximated by a finite one. For example, the Taylor series expansion of functions is often truncated after a few terms, introducing truncation errors.
  • Approximation Error[13, 65]: This occurs when an exact mathematical solution is approximated using a numerical method. For instance, when we approximate a continuous function by a finite sum, an error is introduced.
@@ 22.3 Error Propagation
Error propagation refers to how numerical errors (such as round-off and truncation errors) accumulate and affect the final result of a computation[18]. In deep learning, error propagation is especially important in gradient-based optimization algorithms, where small errors in gradient computations can impact the convergence of the model.
@@ 22.4 Absolute and Relative Error
  • Absolute Error[274, 49, 270, 20]: The difference between the exact value and the approximate value.
    Absolute $\text{Absolute Error}=|x_{\text{exact}}-x_{\text{approximate}}|$
  • Relative Error[37, 187, 290, 43]: The absolute error divided by the exact value, providing a normalized measure of the error.
    Relative $\text{Relative Error}=\frac{|x_{\text{exact}}-x_{\text{approximate}}|}{|x_{\text{exact}}|}$
@@ 22.6 Summary
Root finding is a fundamental problem in numerical analysis and computational mathematics[160]. The objective is to find solutions, or ”roots,” of equations of the form $f(x)=0$. Root finding methods are essential in a wide range of scientific and engineering problems, where an exact algebraic solution is either impossible or impractical to obtain. This chapter introduces several popular methods for finding roots, explains their mathematical foundations, and provides step-by-step implementations in Python.
@@ 23.1 Introduction to Root Finding
For example, finding the roots of the equation $f(x)=x^{2}-4$ involves solving $x^{2}-4=0$, whose solutions are $x=2$ and $x=-2$.
@@ 23.5 Fixed-Point Iteration
Fixed-point iteration is a method for finding a root of the equation $f(x)=0$ by rewriting it in the form $x=g(x)$[198]. The idea is to iteratively apply the function $g(x)$ until convergence is achieved.
@@ 23.6.4 Fixed-Point Iteration
Interpolation and function approximation are fundamental concepts in both mathematics and machine learning[188, 33, 231, 175]. In this chapter, we will explore various methods for interpolating data points and approximating functions, which are widely used in numerical analysis, scientific computing, and deep learning. We will begin with basic interpolation techniques such as polynomial interpolation and then move to more advanced methods like spline interpolation and piecewise linear interpolation. Finally, we will discuss how neural networks are used for function approximation in the context of deep learning.
@@ 24.2 Polynomial Interpolation
Polynomial interpolation is a process where we find a single polynomial $P(x)$ that passes through all the given data points[152, 68, 84, 299]. The degree of the polynomial is determined by the number of points: for $n+1$ points, the interpolating polynomial will have degree $n$.
@@ 24.2.1 Lagrange Interpolation
Lagrange interpolation is one of the simplest methods for polynomial interpolation[57, 156, 104] . It constructs the interpolating polynomial by using the concept of Lagrange basis polynomials.
@@ 24.2.2 Newton’s Divided Difference Interpolation
Newton’s divided difference interpolation is another method for constructing an interpolating polynomial[143]. It uses a recursive process to compute the coefficients of the polynomial based on divided differences of the data points.
@@ 24.3 Spline Interpolation
Spline interpolation uses piecewise polynomials to interpolate data[243, 255, 185, 41]. Unlike high-degree polynomial interpolation, which can suffer from oscillations (known as Runge’s phenomenon), spline interpolation ensures smoothness by using lower-degree polynomials over each subinterval between data points.
@@ 24.4 Piecewise Linear Interpolation
Piecewise linear interpolation connects each pair of data points with a straight line[69, 217, 229, 90, 7]. It is a simple form of interpolation and works well when the data points are close together or if high accuracy is not required. It does not guarantee smoothness, but it is computationally efficient.
@@ 24.5.1 Approximating Nonlinear Functions with Neural Networks
Neural networks can approximate any continuous function given enough neurons and layers, according to the Universal Approximation Theorem[65, 123]. In the context of deep learning, function approximation is critical for tasks such as regression, where the goal is to predict a continuous output from input features.
@@
6class SimpleNN(nn.Module):
@@
9 self.fc1 = nn.Linear(1, 10)
10 self.fc2 = nn.Linear(10, 1)
@@ 25.1.1 Finite Difference Methods
Finite difference methods are the most common numerical techniques for estimating derivatives[99, 128, 34, 121, 129]. They approximate the derivative of a function by considering the differences between function values at discrete points.
@@
The backward difference method approximates the derivative by looking at the difference between the function values at $x$ and a previous point $x-h$:
@@ 25.2 Introduction to Numerical Integration
Numerical integration is used to estimate the value of a definite integral when the analytic solution is difficult or impossible to obtain[54]. Several methods exist for numerical integration, each with varying degrees of accuracy and complexity.
@@ 25.2.1 Trapezoidal Rule
The trapezoidal rule is one of the simplest methods for numerical integration[222, 5]. It approximates the area under a curve by dividing it into trapezoids, calculating the area of each, and summing them up. The integral of a function $f(x)$ over the interval $[a,b]$ is approximated as:
[…]
@@
In this example, we estimated the integral of $\sin(x)$ over $[0,\pi]$ using the trapezoidal rule.

### v2

@@ 15.2 Frobenius Norm
The Frobenius norm is a matrix norm equivalent to the L2 norm for vectors[Recht2010]. It is defined as the square root of the sum of the absolute squares of the matrix elements:
@@ 15.3 Cosine Similarity
Cosine similarity measures the cosine of the angle between two vectors[manning2008introduction]. It is often used in text analysis and other fields where the direction of vectors is more important than their magnitude.
@@
$\text{cosine similarity}=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$
@@ 15.4 Euclidean Distance
Euclidean distance is the straight-line distance between two points in Euclidean space[Curriero2006]. For two vectors $\mathbf{a}$ and $\mathbf{b}$, it is defined as:
@@
Automatic differentiation (AD) is a powerful technique used in many machine learning frameworks, including PyTorch and TensorFlow, to compute gradients efficiently and accurately[baydin2018automatic]. Unlike symbolic differentiation, which can produce complex expressions, or numerical differentiation, which can suffer from precision issues, automatic differentiation computes derivatives systematically using the chain rule. This chapter introduces the concept of automatic differentiation and demonstrates how gradients can be computed in popular machine learning libraries like PyTorch and TensorFlow.
@@ 16.1 Introduction to Automatic Differentiation
Reverse-mode AD is particularly useful in deep learning, where we often need to compute the gradient of a loss function with respect to model parameters[franceschi2017forward].
@@ 16.4 Jacobian and Hessian Computation
In some cases, we may need to compute higher-order derivatives, such as the Jacobian matrix or Hessian matrix[karpfinger2022calculus]. The Jacobian represents the matrix of first-order partial derivatives for a vector-valued function, while the Hessian is the matrix of second-order partial derivatives.
@@ 16.5 Summary
In this chapter, we will introduce the basic concepts behind optimization in deep learning, starting with Gradient Descent, which is the foundation of many advanced optimization techniques. We will then explore Stochastic Gradient Descent (SGD)[robbins1951stochastic], momentum-based optimization, and adaptive optimization methods like Adagrad[duchi2011adaptive], RMSprop[tieleman2012lecture], Adam[kingma2014adam], and AdamW[loshchilov2017decoupled].
@@ 17.1 Gradient Descent
Gradient Descent is one of the most fundamental optimization algorithms used in deep learning. The basic idea is to minimize a loss function (also called the objective function) by moving in the direction of the negative gradient, which points towards the steepest descent[bengio2012practical].
@@ 17.2 Stochastic Gradient Descent (SGD)
While Gradient Descent computes the gradient using the entire dataset, this can be computationally expensive for large datasets. Stochastic Gradient Descent (SGD) solves this problem by approximating the gradient using a small batch of the data, or even a single data point[newton2018recent].
@@ 17.3 Momentum-based Optimization
Momentum-based optimization improves the convergence speed of Gradient Descent by adding momentum to the update rule[cutkosky2019momentum]. This allows the optimizer to continue moving in the same direction if the gradient consistently points in the same direction, avoiding oscillations and speeding up convergence.
@@ 17.4 Adaptive Optimization Methods
Adaptive optimization methods automatically adjust the learning rate during training, which can significantly improve convergence[liu2019variance]. These methods include Adagrad, RMSprop, Adam, and AdamW, each of which modifies the learning rate based on the gradients.
@@ 17.5 Learning Rate Schedules
The learning rate is one of the most important hyperparameters in deep learning. It controls how much the model’s weights are adjusted with respect to the loss gradient during training. Choosing the correct learning rate can significantly impact model performance and training time. However, the optimal learning rate often changes throughout training. This is where learning rate schedules come into play[DarEl2013].
A learning rate schedule adjusts the learning rate dynamically during training, helping the model converge faster and avoid getting stuck in local minima. There are several common strategies for learning rate scheduling, including step decay, exponential decay, and warm restarts[loshchilov2017sgdr].
@@ 17.5.1 Step Decay
Step decay is one of the simplest learning rate schedules. In step decay, the learning rate is reduced by a constant factor at predetermined intervals (epochs)[wang2021convergence]. The idea is that a larger learning rate is beneficial at the start of training, but as the model approaches convergence, reducing the learning rate allows finer adjustments to the weights.
@@ 17.5.2 Exponential Decay
Exponential decay is another commonly used learning rate schedule. Instead of reducing the learning rate in steps, exponential decay reduces the learning rate continuously over time according to an exponential function[lewkowycz2021how]. This allows for a more gradual and smoother reduction in the learning rate.
@@ 18.1 Batch Normalization in Training
Batch normalization is a technique that helps to stabilize and accelerate the training of deep neural networks by normalizing the inputs to each layer[ioffe2015batch]. This process ensures that the input distribution to each layer remains stable during training, making the network less sensitive to the initialization of parameters and improving convergence.
@@ 18.2 Gradient Clipping
Gradient clipping is a technique used to prevent the problem of exploding gradients during the training of deep neural networks[pascanu2013difficulty, liu2021activated]. When gradients become excessively large, they can cause updates to the network weights that are too drastic, leading to unstable training or even causing the model to diverge. Gradient clipping limits the size of the gradients by setting a maximum threshold.
@@ 18.2.1 How Does Gradient Clipping Work?
Given a gradient $g$ and a threshold $t$, if g>t||g||>t, we rescale the gradient:
[…]
@@ 18.3 Second-order Optimization Methods
Second-order optimization methods are more sophisticated than standard first-order methods like gradient descent[battiti1992first]. These methods take into account not only the gradient (first derivative) but also the curvature of the objective function (second derivative). This can lead to faster convergence in some cases, especially when the objective function has complex curvature.
@@ 18.3.1 Newton’s Method
Newton’s Method is a second-order optimization technique that uses both the gradient and the Hessian matrix (the matrix of second-order partial derivatives) to update parameters[powell1974overview]. The basic update rule in Newton’s Method is:
@@ 18.3.2 Quasi-Newton Methods
Quasi-Newton methods are a family of optimization algorithms that approximate the Hessian matrix rather than computing it explicitly[ford1994multi]. This makes them more computationally efficient than full Newton’s Method, especially for high-dimensional problems. One of the most popular Quasi-Newton methods is the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm[goldfarb1970family].
@@ 20.2 Basic Gradient Computation Problems
Gradient computations are essential for optimization in deep learning, as they enable the backpropagation algorithm to update network weights[bengio2012practical]. In this section, we will practice calculating gradients manually and using automatic differentiation tools like those in NumPy.
@@ 20.4 Real-World Linear Algebra Applications in Deep Learning
In deep learning, linear algebra concepts like matrix multiplication, matrix inversion, and eigenvalue decomposition are used extensively[aggarwal2020linear]. Here are some real-world applications:
@@ 22.2 Sources of Errors in Computation
  • Round-off Error[demmel1992round]: This occurs due to the limited precision of floating-point arithmetic used by computers. For example, irrational numbers like $\pi$ and square roots of non-perfect squares cannot be represented exactly.
  • Truncation Error[stoer2013numerical]: This occurs when an infinite process is approximated by a finite one. For example, the Taylor series expansion of functions is often truncated after a few terms, introducing truncation errors.
  • Approximation Error[bartlett2005bound, cybenko1989approximation]: This occurs when an exact mathematical solution is approximated using a numerical method. For instance, when we approximate a continuous function by a finite sum, an error is introduced.
@@ 22.3 Error Propagation
Error propagation refers to how numerical errors (such as round-off and truncation errors) accumulate and affect the final result of a computation[bevington2002data]. In deep learning, error propagation is especially important in gradient-based optimization algorithms, where small errors in gradient computations can impact the convergence of the model.
@@ 22.4 Absolute and Relative Error
  • Absolute Error[vapnik1995nature, chen2004measure, valiant1984theory, bishop2006pattern]: The difference between the exact value and the approximate value.
    $\text{Absolute Error}=|x_{\text{exact}}-x_{\text{approximate}}|$
  • Relative Error[burden2016numerical, mclean2010introduction, wrench1963relative, chapra2010numerical]: The absolute error divided by the exact value, providing a normalized measure of the error.
    $\text{Relative Error}=\frac{|x_{\text{exact}}-x_{\text{approximate}}|}{|x_{\text{exact}}|}$
@@ 22.6 Summary
Root finding is a fundamental problem in numerical analysis and computational mathematics[leader2022numerical]. The objective is to find solutions, or ”roots,” of equations of the form $f(x)=0$. Root finding methods are essential in a wide range of scientific and engineering problems, where an exact algebraic solution is either impossible or impractical to obtain. This chapter introduces several popular methods for finding roots, explains their mathematical foundations, and provides step-by-step implementations in Python.
@@ 23.1 Introduction to Root Finding
For example, finding the roots of the equation $f(x)=x^{2}-4$ involves solving $x^{2}-4=0$, whose solutions are $x=2$ and x=2x=-2.
@@
  • The Bisection Method[press2007numerical, golub1993introduction, chapra2010numerical, burden2016numerical]
  • Newton’s Method[burden2015numerical, mann1943introduction, encyclopediaNewtonRaphson, lancaster1956newton, watson1944treatise]
  • The Secant Method[rice1960secant, burden2016numerical, cheney2009introduction, chapra2016numerical]
  • Fixed-Point Iteration[burden2001first, burden2016numerical, rennick1968fixed, chapra2011numerical, karris2011numerical]
@@ 23.5 Fixed-Point Iteration
Fixed-point iteration is a method for finding a root of the equation $f(x)=0$ by rewriting it in the form x=g(x)x=g(x)[nevanlinna2012convergence]. The idea is to iteratively apply the function $g(x)$ until convergence is achieved.
@@ 23.6.4 Fixed-Point Iteration
Interpolation and function approximation are fundamental concepts in both mathematics and machine learning[micchelli1980function, burden2011numerical, rivlin2003interpolation, lorentz1966approximation]. In this chapter, we will explore various methods for interpolating data points and approximating functions, which are widely used in numerical analysis, scientific computing, and deep learning. We will begin with basic interpolation techniques such as polynomial interpolation and then move to more advanced methods like spline interpolation and piecewise linear interpolation. Finally, we will discuss how neural networks are used for function approximation in the context of deep learning.
@@ 24.2 Polynomial Interpolation
Polynomial interpolation is a process where we find a single polynomial $P(x)$ that passes through all the given data points[kozachenko1992polynomial, davis1967newton, fleming1977interpolation, zayed2011numerical]. The degree of the polynomial is determined by the number of points: for $n+1$ points, the interpolating polynomial will have degree $n$.
@@ 24.2.1 Lagrange Interpolation
Lagrange interpolation is one of the simplest methods for polynomial interpolation[colquhoun1997numerical, lagrange1859new, golub1993introduction] . It constructs the interpolating polynomial by using the concept of Lagrange basis polynomials.
@@ 24.2.2 Newton’s Divided Difference Interpolation
Newton’s divided difference interpolation is another method for constructing an interpolating polynomial[kellogg1975numerical]. It uses a recursive process to compute the coefficients of the polynomial based on divided differences of the data points.
@@ 24.3 Spline Interpolation
Spline interpolation uses piecewise polynomials to interpolate data[schoenberg1946contributions, stoer2013introduction, mazure2001spline, carroll2006interpolation]. Unlike high-degree polynomial interpolation, which can suffer from oscillations (known as Runge’s phenomenon), spline interpolation ensures smoothness by using lower-degree polynomials over each subinterval between data points.
@@ 24.4 Piecewise Linear Interpolation
Piecewise linear interpolation connects each pair of data points with a straight line[deboor1972calculating, powell1981piecewise, rice1960secant, friedman1984proof, atkinson1989introduction]. It is a simple form of interpolation and works well when the data points are close together or if high accuracy is not required. It does not guarantee smoothness, but it is computationally efficient.
@@ 24.5.1 Approximating Nonlinear Functions with Neural Networks
Neural networks can approximate any continuous function given enough neurons and layers, according to the Universal Approximation Theorem[cybenko1989approximation, hornik1989multilayer]. In the context of deep learning, function approximation is critical for tasks such as regression, where the goal is to predict a continuous output from input features.
@@
6class SimpleNN($n$.Module):
@@
9 self.fc1 = $n$.Linear(1, 10)
10 self.fc2 = $n$.Linear(10, 1)
@@ 25.1.1 Finite Difference Methods
Finite difference methods are the most common numerical techniques for estimating derivatives[gibson1980finite, ince1956numerical, burden2016finite, hindmarsh1973finite, incropera2002numerical]. They approximate the derivative of a function by considering the differences between function values at discrete points.
@@
The backward difference method approximates the derivative by looking at the difference between the function values at $x$ and a previous point xhx-h:
@@ 25.2 Introduction to Numerical Integration
Numerical integration is used to estimate the value of a definite integral when the analytic solution is difficult or impossible to obtain[ciarlet2002handbook]. Several methods exist for numerical integration, each with varying degrees of accuracy and complexity.
@@ 25.2.1 Trapezoidal Rule
The trapezoidal rule is one of the simplest methods for numerical integration[press2007numerical, ahlberg1968methods]. It approximates the area under a curve by dividing it into trapezoids, calculating the area of each, and summing them up. The integral of a function $f(x)$ over the interval $[a,b]$ is approximated as:
$\int_{a}^{b}f(x)\,dx\approx\frac{h}{2}\left[f(a)+2\sum_{i=1}^{n-1}f(x_{i})+f(b)\right]$
@@
In this example, we estimated the integral of $\sin(x)$ over $0,\pi]$ using the trapezoidal rule.